David signed a contract to build 40 bird houses for the local pet shop. The owner wants them ready within the month. For every one David produces he will receive $20.00 and for every one he fails to produce he will be fined $8.00. At the end of the month David received $380.00. How many bird houses did he build?
Answer: 25 (How?)

Answer by fractalier(6550)

Call the number he completed, x. Then the number he didn't complete would be 40-x. Do you see that? If you do, we then write
20x - 8(40-x) = 380
20x - 320 + 8x = 380
28x - 320 = 380
28x = 700
x = 700/28 = 25 bird houses he completed